# Flow Matching for Tabular Data Synthesis

### v1

@@ D.1 Variance Comparison Results
Table 8: Comparative performance of synthetic data results using theoretical (left) and relaxed (right) variance for TabbyFlow-OT. The table highlights that relaxing the theoretical variance to learn the distribution of continuous variables frequently leads to better performance.
@@
D.2 Main Comparison on ID, AD, CH Datasets
Table 9 presents a comparative evaluation of six synthetic data generation algorithms across four datasets (UK, ID, AD, CH), highlighting the two best-performing algorithms per dataset. Overall, TabSynFlow-OT consistently achieves superior utility in the ID and CH datasets compared to diffusion baselines, and matches TabSyn while delivering lower disclosure risk in AD. Notably, TabSynFlow-OT also provides reduced risk in ID and CH, making it an effective choice for synthesising tabular data which increases utility and reduces privacy risk.
By contrast, TabbyFlow-OT yields the highest utility for AD, a result that is accompanied by increased disclosure risk, while in ID, it generates synthetic data with lower utility but achieves a substantial reduction in risk. This pattern illustrates how variations in model performance and privacy exposure reflect the empirical relationship between utility and risk across different datasets. TabSyn consistently ranks as the second-best in utility across all three datasets.
By combining with the main text, these results underscore the strength of flow matching approaches for tabular data synthesis, particularly the flexibility of TabSynFlow-OT as an alternative to diffusion-based models in extreme number of observations, i.e. too high or too low.
Table 9: Comparative Performance of six synthetic data generation algorithms across three datasets (ID, AD, CH). Bold and underline indicate the first and second best performing algorithms for each dataset, respectively. The table highlights the comparable performance of TabbyFlow and TabSyn, while TabSynFlow-OT stands out as a strong alternative—particularly in the ID dataset, where TabbyFlow exhibits higher risk in conjunction with generated high utility synthetic data.
Evaluation Algorithm ID AD CH
@@ D.2 Main Comparison on ID, AD, CH Datasets
D.3 Number of Function Evaluations
Figure 3 shows how computational cost is related to the quality of the output between methods (TabSyn, TabSynFlow and TabbyFlow) and datasets. In general, both utility and risk increase as NFEs grow. In four datasets (UK, RW, CA, FI), TabbyFlow seems to outperform TabSyn. However, in ID and CH, TabSynFlow-OT outperforms TabSyn. Therefore, if one has a limited computational budget, they may choose TabbyFlow or TabSynFlow using the OT path as the algorithm that performed best on low-step setting among all algorithms, particularly TabSyn.
D.4 Integration Time
While the original implementation of flow matching uses full ODE integration, it is possible to terminate the ODE at intermediate times—a principle explored in this section. Our analysis of ODE integration times ($t_{ode}$) in Figure 4, spanning datasets including Rwanda, Adult, Churn, UK, Indonesia, Canada, and Fiji, offers critical insights into how utility and risk evolve during synthetic data generation. For TabSynFlow with optimal transport (OT), early stopping at $t_{ode}=0.6$ consistently produces high-utility synthetic data, whereas variance preserving (VP) paths in both TabSynFlow and TabbyFlow tend to underperform at this stage. This indicates that OT trajectories better retain the data structure in earlier integration steps, while VP introduces noise prematurely. Notably, the Rwanda dataset presents an interesting anomaly: TabbyFlow-OT achieves slightly better utility than VP methods at $t_{ode}=0.6$, which may be attributed to the dataset’s relatively simple feature distributions.
Figure 5 showed the utility and risk in details when $0.9\leq t_{ode}\leq 1$. At $t=0.9$, OT methods maintain superior performance, particularly for the UK and Adult datasets, although the gap between OT and VP narrows for the latter. This implies dataset-specific sensitivity to VP’s noise characteristics. Near full integration ($t_{ode}\rightarrow 1.0$), all algorithms exhibit erratic behaviour in some datasets such as Fiji: the utility drops and the risk increases. This problem suggests that flow matching is prone to quality drops when fully integrated.
Figure 4: Effect of ODE integration time on utility–risk (per dataset + average) for TabSynFlow and TabbyFlow. Utility and risk evaluation of TabSynFlow and TabbyFlow based on dataset and average on different integration time from $t_{ode}=[0.6,1]$ with interval 0.1. Colors encode method and path (VP vs OT). OT achieves high-utility solutions early (e.g., $t_{ode}=0.6$), while pushing to $t_{ode}\to 1$ often reduces utility and/or increases risk—so early stopping is preferable.
Figure 5: Utility and risk evaluation of TabSynFlow and TabbyFlow based on dataset and average on late integration time $t=[0.9,0.95,0.975,1]$. For selected datasets, we zoom into the late integration window using the same color/marker scheme as Figure 4. Take-home: Near full integration, instability emerges across several datasets (e.g., Fiji), consistent with utility drops and risk increases; OT remains comparatively robust but still degrades as $t_{ode}\to 1$.
Appendix E Flow Matching Results on Different Interpolants
E.1 Interpolation Path in Flow Matching
A central element in continuous-time generative modelling is the choice of interpolation schedule, which determines how the clean data $x_{1}$ is progressively transformed into noise, and conversely how the noise is mapped back to the data. This schedule is characterised by two functions: the signal decay factor $\alpha(t)$ and the noise scale $\sigma(t)$. Together, these define the marginal distribution.
$x_{t}\sim\mathcal{N}\big(\alpha(t)x_{1},\sigma^{2}(t)I\big),$
Different schedules entail distinct trade-offs between training stability, sample quality, and computational efficiency. While the main text provides the primer for the OT and VP paths, flow matching formulations can also use alternative paths such as variance-exploding (lipman2023flowmatchinggenerativemodeling), cosine (albergo2023buildingnormalizingflowsstochastic), or logit-normal (labs2025flux1kontextflowmatching) interpolants. In this section, we summarise these schedules in two comparative tables: Table 10 for the interpolants $(\alpha(t),\sigma(t))$ and Table 11 for their derivatives $(\dot{\alpha}(t),\dot{\sigma}(t))$.
Interpolation
$\alpha(t)$ $\sigma(t)$
Linear 1t1-t $t$
Cosine $\sin\!\big(\dfrac{\pi}{2}t\big)$ $\cos\!\big(\dfrac{\pi}{2}t\big)$
Logit-Normal
$\dfrac{e^{\mu}}{e^{\mu}+(\tfrac{1}{ct}-1)^{\lambda}}$
$\mu=\log 3,\,\lambda=1$
$1-\alpha(t)$
Table 11: Derivative of interpolants for different schedules.
Interpolation
$\dot{\alpha}(t)$ $\dot{\sigma}(t)$
Linear 1-1 11
E.2 Empirical Results on TabSynFlow
Dataset OT VP VE Cosine Logit-Normal
Utility Risk Utility Risk Utility Risk Utility Risk Utility Risk
E.3 Empirical Results on TabbyFlow
Table 13: Comparison of Utility and Risk Across Different Paths on TabbyFlow
Dataset OT VP VE Cosine Logit-Normal
Utility Risk Utility Risk Utility Risk Utility Risk Utility Risk
Indonesia 0.6671 0.4114 0.6488 0.3918 0.4519 0.3955 0.6319 0.3856 0.6520 0.3888
E.4 SDE Performance on Different $g_{t}$
Tables 14 and 15 disseminate the comparison of the difference in utility ($\Delta\text{U}$) and risk ($\Delta\text{R}$) when using SDE with different $g_{t}$ on TabbyFlow compared to its ODE counterparts as a baseline (See Table 13). The tables reveal consistently small deviations across all datasets, with the majority of differences within [0.02,+0.02][-0.02,+0.02]. This minimal variation strongly validates the theoretical equivalence between the SDE and ODE marginals, demonstrating that SDEs preserve the same distributional properties regardless of the specific $g_{t}$ function employed. The results confirm that the controlled stochasticity introduced through SDE sampling maintains distributional fidelity while offering additional flexibility in the sampling process.
Table 14: Difference of utility and risk across different $g_{t}$ setting in SDE w.r.t. the ODE results on TabbyFlow-OT.
Dataset OT VP VE Cos Logit
Table 15: Difference of utility and risk across different $g_{t}$ setting in SDE w.r.t. the ODE results on TabbyFlow-VP.
Dataset OT VP VE Cos Logit

### v2

@@ E.1 Results on Standard Tabular Generative Model Evaluations
[…]
Mixed-type pairwise fidelity (Trend).
Trend evaluates how well pairwise dependence is preserved, combining correlation discrepancies for continuous pairs and contingency-table discrepancies for categorical pairs. Table 12 largely mirrors the Shape findings. TabbyFlow achieves the lowest Trend error on UK, Canada, Fiji, and Rwanda, indicating improved preservation of pairwise structure on these datasets. However, TabSynFlow-OT is best on Indonesia, Adult, and Churn, and TabbyFlow performs poorly on Churn in particular, where its Trend error is several times larger than competing methods. Overall, these results suggest that TabbyFlow’s dependence preservation is dataset-dependent, and its strongest gains occur on the census-style datasets with richer categorical structure, while TabSynFlow-OT provides more stable pairwise behavior on Adult, Churn, and Indonesia.
Table 12: Performance comparison on the error rates (%) of Trend. We report the mean $\pm$ standard deviation across 20 seeds of synthetic data generation. Bold and underline indicate the first and second best performing algorithms for each dataset, respectively. The table highlights the domination of TabbyFlow and TabSynFlow as the best performers, meaning that the algorithms are best in preserving the correlation and contingency of mixed-type data.
Data TabDDPM TabSyn TabSynFlow TabSynFlow TabbyFlow TabbyFlow
OT VP OT VP
[…]
Detectability (C2ST).
Table 13 reports C2ST detection scores using logistic regression. Across most datasets, TabbyFlow variants achieve the strongest detection scores, consistent with producing samples that are harder to distinguish from real data under the chosen detector. There are notable exceptions. On Indonesia, TabSynFlow-OT attains the top score, and on Churn, TabSyn and TabSynFlow-OT outperform TabbyFlow. TabDDPM is highly inconsistent across datasets, with extremely low scores on Canada and Fiji and much higher scores elsewhere, indicating substantial dataset sensitivity for this detector-based criterion.
Table 13: Performance comparison on the detection score (C2ST) using logistic regression classifier. . We report the mean $\pm$ standard deviation across 20 seeds of synthetic data generation. Bold and underline indicate the first and second best performing algorithms for each dataset, respectively. The table highlights the domination of TabbyFlow as the best performers, meaning that the algorithms are best in producing realistic synthetic data.
Data TabDDPM TabSyn TabSynFlow TabSynFlow TabbyFlow TabbyFlow
OT VP OT VP
UK 0.9563 0.9561 0.9652 0.8801 0.9789 0.9827
Canada 0.0555 0.8439 0.8404 0.8530 0.8770 0.8712
Fiji 0.0376 0.9763 0.9815 0.9550 0.9821 0.9859
Rwanda 0.4347 0.9824 0.9779 0.9632 0.9932 0.9972
Indonesia 0.9920 0.9891 0.9983 0.9887 0.9898 0.9808
Adult 0.9427 0.9704 0.9826 0.9458 0.9639 0.9881
Churn 0.9748 0.9957 0.9949 0.8700 0.9222 0.9270
Sample quality and coverage ($\alpha$-precision and $\beta$-recall).
Tables 1415 report $\alpha$-precision (sample fidelity) and $\beta$-recall (coverage) averaged over 20 generation seeds. Across all datasets, $\alpha$-precision is high for TabSynFlow and TabbyFlow, indicating that their synthetic records typically fall within high-density regions of the real data. TabSynFlow is best on four datasets (Fiji, Rwanda, Adult, and Churn) and second-best on two (UK and Indonesia), while TabbyFlow is best on three datasets (UK, Canada, and Indonesia) and frequently appears among the top two, showing that both families produce high-fidelity samples.
In contrast, $\beta$-recall reveals clearer dataset dependence. TabbyFlow achieves the best coverage on five datasets (UK, Canada, Fiji, Rwanda, and Adult), suggesting that it more consistently captures diverse modes of the real distribution rather than concentrating on a narrow high-density region. However, Indonesia and Churn are exceptions: on Indonesia, TabSyn and TabSynFlow are slightly higher than TabbyFlow, and on Churn, TabDDPM achieves the highest $\beta$-recall. Overall, these results indicate that TabSynFlow and TabbyFlow are comparably strong in sample-level fidelity, while TabbyFlow more often improves coverage, with a small number of datasets where alternative methods provide better recall.
Table 14: Performance comparison on $\alpha$-precision (higher is better). We report mean $\pm$ standard deviation across 20 synthetic generation seeds. Bold and underline indicate the best and second-best methods per dataset. Overall, TabSynFlow and TabbyFlow consistently achieve the highest $\alpha$-precision, indicating strong sample-level fidelity across datasets.
Data TabDDPM TabSyn TabSynFlow TabSynFlow TabbyFlow TabbyFlow
OT VP OT VP
[…]
Table 15: Performance comparison on $\beta$-recall (higher is better). We report mean $\pm$ standard deviation across 20 synthetic generation seeds. Bold and underline indicate the best and second-best methods per dataset. TabbyFlow attains the highest $\beta$-recall on most datasets, indicating stronger coverage of the real data distribution
Data TabDDPM TabSyn TabSynFlow TabSynFlow TabbyFlow TabbyFlow
OT VP OT VP
[…]
E.2 Number of Function Evaluations
Figure 3 shows how computational cost is related to the quality of the output between methods (TabSyn, TabSynFlow and TabbyFlow) and datasets. In general, both utility and risk increase as NFEs grow. In four datasets (UK, RW, CA, FI, ID), TabbyFlow seems to outperform TabSyn. However, CH, TabSynFlow-OT outperforms TabSyn. Therefore, if one has a limited computational budget, they may choose TabbyFlow or TabSynFlow using the OT path as the algorithm that performed best on low-step setting among all algorithms, particularly TabSyn.
E.3 Integration Time
While the original implementation of flow matching uses full ODE integration, it is possible to terminate the ODE at intermediate times—a principle explored in this section. Our analysis of ODE integration times ($t_{\text{ode}}$) in Figure 4, spanning datasets including Rwanda, Adult, Churn, UK, Indonesia, Canada, and Fiji, offers critical insights into how utility and risk evolve during synthetic data generation. For TabSynFlow with optimal transport (OT), early stopping at $t_{\text{ode}}=0.6$ consistently produces high-utility synthetic data, whereas variance preserving (VP) paths in both TabSynFlow and TabbyFlow tend to underperform at this stage. This indicates that OT trajectories better retain the data structure in earlier integration steps, while VP introduces noise prematurely. Notably, the Rwanda dataset presents an interesting anomaly: TabbyFlow-OT achieves slightly better utility than VP methods at $t_{\text{ode}}=0.6$, which may be attributed to the dataset’s relatively simple feature distributions.
Figure 5 showed the utility and risk in details when $0.9\leq t_{\text{ode}}\leq 1$. At $t=0.9$, OT methods maintain superior performance, particularly for the UK and Adult datasets, although the gap between OT and VP narrows for the latter. This implies dataset-specific sensitivity to VP’s noise characteristics. Near full integration ($t_{\text{ode}}\rightarrow 1.0$), all algorithms exhibit erratic behaviour in some datasets such as Fiji: the utility drops and the risk increases. This problem suggests that flow matching is prone to quality drops when fully integrated.
Figure 4: Effect of ODE integration time on utility–risk (per dataset + average) for TabSynFlow and TabbyFlow. Utility and risk evaluation of TabSynFlow and TabbyFlow based on dataset and average on different integration time from $t_{\text{ode}}=[0.6,1]$ with interval 0.1. Colors encode method and path (VP vs OT). OT achieves high-utility solutions early (e.g., $t_{\text{ode}}=0.6$), while pushing to $t_{\text{ode}}\to 1$ often reduces utility and/or increases risk—so early stopping is preferable.
Figure 5: Utility and risk evaluation of TabSynFlow and TabbyFlow based on dataset and average on late integration time $t=[0.9,0.95,0.975,1]$. For selected datasets, we zoom into the late integration window using the same color/marker scheme as Figure 4. Take-home: Near full integration, instability emerges across several datasets (e.g., Fiji), consistent with utility drops and risk increases; OT remains comparatively robust but still degrades as $t_{\text{ode}}\to 1$ in some cases.
Appendix F Flow Matching Results on Different Interpolants
F.1 Interpolation Path in Flow Matching
A central element in continuous-time generative modelling is the choice of interpolation schedule, which determines how the clean data $x_{1}$ is progressively transformed into noise, and conversely how the noise is mapped back to the data. This schedule is characterised by two functions: the signal decay factor $\alpha(t)$ and the noise scale $\sigma(t)$. Together, these define the conditional distribution.
$x_{t}|x_{1}\sim\mathcal{N}\big(\alpha(t)x_{1},\sigma^{2}(t)I\big),$
F.2 Trajectory Setting in Flow Matching
There are two commonly used trajectories to define the probability path in flow matching.
The first is the optimal transport (OT) trajectory (Liu et al. 2022; Lipman et al. 2023), where the mean and standard deviation of the conditional distribution evolve linearly over time, resulting in straight-line (conditional) paths connecting the noise and data distributions. OT-based trajectories are computationally straightforward, efficient, and easy to parameterise due to their time-invariant structure, as shown in equation F.1
$\begin{split}x_{t}&=tx_{1}+(1-(1-\sigma_{\text{min}})t)x_{0}\\ u_{t}(x_{t}\mid x_{1})&=\frac{x_{1}-(1-\sigma_{\text{min}})x_{t}}{1-(1-\sigma_{\text{min}})t},\end{split}$ (F.1)
The second type of trajectory is the variance-preserving (VP) path, derived from stochastic differential equations commonly used in diffusion models (Lipman et al. 2023). In this approach, noise is gradually added to the data during the forward process, while the reverse process learns to de-noise using time-dependent vector fields. In contrast to the straight OT path, where the conditional mean evolves linearly in $t$, the VP path follows a noise schedule that induces non-uniform change over time, with smaller updates in early times and larger updates later. The parameters $\alpha_{t}$ and $\sigma_{t}$ are determined by cumulative functions of the noise scale function $\beta(t)=\beta_{\text{min}}+t(\beta_{\text{max}}-\beta_{\text{min}})$, as given in equation F.2.
$\begin{split}\alpha_{t}&=e^{-\frac{1}{2}T(t)},\quad T(t)=\int_{0}^{1-t}\beta(s)ds,\quad\sigma_{t}=\sqrt{1-\alpha_{t}^{2}}\\ u_{t}(x_{t}|x_{1})&=-\frac{\dot{T}(t)}{2}\left[\frac{e^{-T(t)}x_{t}-e^{-\frac{1}{2}T(t)}x_{1}}{1-e^{-T(t)}}\right]\end{split}$ (F.2)
The choice of trajectory – or interpolation scheme – significantly affects data synthesis performance, which we explore in this paper. Depending on the data structure or modelling objectives, either trajectory can be more suitable. Therefore, flow matching formulations can also use alternative paths such as variance-exploding (Lipman et al. 2023), cosine (Albergo and Vanden-Eijnden 2023), or logit-normal (Black Forest Labs 2025) interpolants. In this section, we summarise these schedules in two comparative tables: Table 16 for the interpolants $(\alpha(t),\sigma(t))$ and Table 17 for their derivatives $(\dot{\alpha}(t),\dot{\sigma}(t))$.
Interpolation $\alpha(t)$ $\sigma(t)$
Linear 1t1-t $t$
Cosine $\sin\!\big(\dfrac{\pi}{2}t\big)$ $\cos\!\big(\dfrac{\pi}{2}t\big)$
Logit-Normal $\dfrac{e^{\mu}}{e^{\mu}+(\tfrac{1}{ct}-1)^{\lambda}}$ $\mu=\log 3,\,\lambda=1$ $1-\alpha(t)$
Table 17: Derivative of interpolants for different schedules.
Interpolation $\dot{\alpha}(t)$ $\dot{\sigma}(t)$
Linear 1-1 11
F.3 Empirical Results on TabSynFlow
Dataset OT VP VE Cosine Logit-Normal
Utility Risk Utility Risk Utility Risk Utility Risk Utility Risk
F.4 Empirical Results on TabbyFlow
Table 19: Comparison of Utility and Risk Across Different Paths on TabbyFlow. We report the mean across 20 seeds of synthetic data generation.
Dataset OT VP VE Cosine Logit-Normal
Utility Risk Utility Risk Utility Risk Utility Risk Utility Risk
Indonesia 0.9191 0.6556 0.8473 0.6447 0.5309 0.5922 0.7775 0.6338 0.8984 0.6365
F.5 SDE Performance on Different $g_{t}$
Tables 20 and 21 disseminate the comparison of the difference in utility ($\Delta\text{U}$) and risk ($\Delta\text{R}$) when using SDE with different $g_{t}$ on TabbyFlow compared to its ODE counterparts as a baseline (See Table 19). The tables reveal consistently small deviations across all datasets, with the majority of differences within [0.02,+0.02][-0.02,+0.02]. This minimal variation strongly validates the theoretical equivalence between the SDE and ODE marginals, demonstrating that SDEs preserve the same distributional properties regardless of the specific $g_{t}$ function employed. The results confirm that the controlled stochasticity introduced through SDE sampling maintains distributional fidelity while offering additional flexibility in the sampling process.
Table 20: Difference of utility and risk across different $g_{t}$ setting in SDE w.r.t. the ODE results on TabbyFlow-OT. We report the mean across 20 seeds of synthetic data generation.
Dataset OT VP VE Cos Logit
Table 21: Difference of utility and risk across different $g_{t}$ setting in SDE w.r.t. the ODE results on TabbyFlow-VP. We report the mean across 20 seeds of synthetic data generation.
Dataset OT VP VE Cos Logit
Appendix G Ablation Studies
G.1 Variance Comparison Results
[…]
@@ G.1 Variance Comparison Results
Table 22: Comparative performance of synthetic data results using theoretical (left) and relaxed (right) variance for TabbyFlow-OT. We report the mean $\pm$ standard deviation across 20 seeds of synthetic data generation. The table highlights that relaxing the theoretical variance to learn the distribution of continuous variables frequently leads to better performance.
@@
G.2 Stress-test: Latent-space vs. Data-space Flow Matching.
G.2.1 High Cardinality Data: Diabetes
Table 23 reports an ablation on the Diabetes dataset ($D_{\text{num}}=8$, $\max(K_{d})=790$, $\sum K_{d}=2513$ [^3] Data can be accessed in https://archive.ics.uci.edu/dataset/296/diabetes+130-us+hospitals+for+years+1999-2008 comparing latent-space flow matching (TabSynFlow) and data-space flow matching (TabbyFlow). Due to the memory constraint in high dimensional setting, we excluded the high cardinality columns when calculating $\alpha$-precision, $\beta$-recall, and WD, and also do batch-split on $\alpha$-precision, $\beta$-recall.
Based on the table, TabbyFlow is stronger at preserving data utility and fidelity, with lower WD, shape, and trend values and higher $\alpha$-precision and $\beta$-recall. In contrast, TabSynFlow better matches continuous marginals, achieving substantially lower WD, and TabSynFlow-OT and TabbyFlow-VP attains the most favorable utility-risk balance among the four configurations. Overall, this stress test suggests a clear division of strengths: with the availability of high category data, data-space VFM (TabbyFlow) improves utility and fidelity, while latent-space FM (TabSynFlow) improves continuous alignment.
Table 23: High-cardinality stress test on the Diabetes dataset ($n=101,766$, $D_{\text{num}}=8$, $\max(K_{d})=790$, $\sum K_{d}=2513$): comparison of latent-space flow matching (TabSynFlow) and data-space flow matching (TabbyFlow) under OT and VP trajectories. We report the mean $\pm$ standard deviation across 20 seeds of synthetic data generation.. The results showed that TabbyFlow performed better in both OT and VP, while in latent space, TabSynFlow-OT showed best performance.
Evaluation TabSynFlow-OT TabSynFlow-VP TabbyFlow-OT TabbyFlow-VP
[…]
WD 2.5976±0.0025 2.7618±0.0026 2.3773±0.0028 2.4788±0.0124
Shape 3.33±0.03 5.92±0.03 1.93±0.03 1.45±0.04
Trend 5.10±0.10 8.85±0.09 6.13±0.18 5.54±0.18
$\alpha$-precision 93.77±0.13 88.2±0.14 93.85±0.15 97.72±0.14
$\beta$-recall 37.01±0.17 27.82±0.14 50.06±0.17 45.81±0.18
G.2.2 High Dimensional Data: Gene Expression
[…]
Table 24: High-cardinality dimensional stress test on the dataset ($D_{\text{num}}=20264$, $D_{\text{cat}}=1$): comparison of latent-space flow matching (TabSynFlow) and data-space flow matching (TabbyFlow) under OT and VP trajectories. We report the mean $\pm$ standard deviation across 20 seeds of synthetic data generation. The results show that TabSynFlow substantially outperforms TabbyFlow on Utility, and Shape in this ultra-high-dimensional regime, while TabbyFlow achieves lower WD over TabSynFlow. Overall, latent-space FM provides a more favorable utility trade-off at >20K dimensions, even with a 256-d projection to avoid OOM, while TabbyFlow preserves joint distribution.
Evaluation TabSynFlow-OT TabSynFlow-VP TabbyFlow-OT TabbyFlow-VP
[…]
WD 212.6±1.04 206.9±1.17 149.8±0.71 155.36±0.62
Shape 9.90±0.08 9.22±0.1 20.21±0.21 20.47±0.18
G.3 Architecture Comparison of TabbyFlow: MLP vs. Transformers
[…]
Table 25: Performance of TabbyFlow when using MLP and transformers architecture. Note that in here we did not use gradual decrease Gaussian loss weight like implemented by Guzmán-Cordero et al. (2025), following our default implementation. We report the mean $\pm$ standard deviation across 20 generation seeds. The results highlights the competitiveness of MLP compared to Transformers in producing synthetic data.
Evaluation Data MLP-OT MLP-VP Transformers-OT Transformers-VP
@@ G.2.2 High Dimensional Data: Gene Expression
[…]